A military helicopter on a training mission is flying horizontally at a speed of 60.0 m/s and accidentally drops a bomb (fortunately not armed) at an elevation of 300 m. You can ignore air resistance.

(a) How much time is required for the bomb to reach the earth?

(b) How far does it travel horizontally while falling?

(c) Find the horizontal and vertical components of its velocity just before it strikes the earth.

(d) Draw x-t, y-t, vx-t, and vy-t graphs for the bomb's motion.

(e) If the velocity of the helicopter remains constant, where is the helicopter when the bomb hits


1
Expert's answer
2021-09-08T16:13:30-0400

(a) "s=gt^2/2\to t=\sqrt{2s/g}=\sqrt{2\cdot300/9.8}=7.82\ (s)"


(b) "l=v\cdot t=60\cdot7.82=469.5\ (m)"


(c) "v_h=v=60\ (m/s)" and "v_v=\sqrt{2sg}=\sqrt{2\cdot300\cdot 9.8}=76.7\ (m/s)"


(d)





(e) "L=v\cdot t=300\cdot7.82=469.2\ (m)" The helicopter will be over the place where the bomb will hit the ground.